# Rainbow subgraphs of uniformly coloured randomly perturbed graphs

### v2

@@ 1 Introduction
Given $\delta\in(0,1)$, we define $\mathcal{G}_{\delta,n}$ to be the family of graphs on vertex set $[n]$ with minimum degree at least $\delta n$, and we let $\mathbf{G}(n,p)$ be the binomial random graph on vertex set $[n]$ with edge probability pp. One of the central themes in extremal combinatorics is determining the minimum degree threshold for a given graph property $\mathcal{P}$, i.e. how large $\delta$ needs to be so that every $G\in\mathcal{G}_{\delta,n}$ satisfies $\mathcal{P}$. Similarly, probabilistic combinatorics aims to determine how large $p$ needs to be for $\mathbf{G}(n,p)$ to satisfy $\mathcal{P}$ with high probability [^1] Formally, we say that a sequence of events $(A_{n})_{n\in\mathbb{N}}$ holds with high probability if $\mathbb{P}[A_{n}]\rightarrow 1$ as $n\rightarrow\infty$. . Bohman, Frieze and Martin [8] provided a connection between the extremal and the random graph settings by introducing the randomly perturbed graph model. For a given $\delta\in(0,1)$, this is defined as $G_{0}\cup\mathbf{G}(n,p)$ where $G_{0}\in\mathcal{G}_{\delta,n}$, i.e. as the graph on $[n]$ whose edge set is the union of the edges of a deterministic graph $G_{0}$ with minimum degree at least $\delta n$ and the edges of a random graph $\mathbf{G}(n,p)$ on the same vertex set. For a given $\delta$ and a given graph property $\mathcal{P}$, a pivotal question in the area is to determine how large $p$ needs to be so that for every $G_{0}\in\mathcal{G}_{\delta,n}$, with high probability, $G_{0}\cup\mathbf{G}(n,p)$ satisfies $\mathcal{P}$. More precisely, we say that $\hat{p}=\hat{p}(\delta,\mathcal{P},n)$ is a perturbed threshold for the property $\mathcal{P}$ at $\delta$ if there are constants $C>c>0$ such that for any $p\geq C\hat{p}$ and for any sequence of $n$-vertex graphs $(G_{n})_{n\in\mathbb{N}}$ with $G_{n}\in\mathcal{G}_{\delta,n}$ we have $\lim_{n\to\infty}\mathbb{P}\big(G_{n}\cup G(n,p)\in\mathcal{P}\big)=1$, and for any $p\leq c\hat{p}$ there exists a sequence of $n$-vertex graphs $(G_{n})_{n\in\mathbb{N}}$ with $G_{n}\in\mathcal{G}_{\delta,n}$ such that $\lim_{n\to\infty}\mathbb{P}\big(G_{n}\cup G(n,p)\in\mathcal{P}\big)=0$.
@@
Given a graph $G$, a vertex $v\in V(G)$ and a subset $X\subseteq V(G)$, $E_{G}(v,X)$ denotes the set of edges of the form vxvx with $x\in X$, and $N_{G}(X)$ denotes the set of vertices with at least one neighbour in $X$. A bare path in $G$ is a path whose interior vertices have degree 22 in $G$. For a graph $G$ and $p\in[0,1]$, the $p$-random subgraph of $G$, denoted by $G_{p}$, is the random graph resulting from sampling each edge of $G$ independently with probability pp.
A digraph $D$ is a set of vertices together with a set of ordered pairs of distinct vertices and the minimum semi-degree $\delta^{0}(D)$ of $D$ is the minimum over in- and out-degrees of vertices in $D$. Moreover $\mathbf{D}(n,p)$ denotes the binomial random digraph on $n$ vertices, that is the digraph on $[n]$ where each ordered pair of distinct vertices forms a directed edge independently with probability pp.
@@ 2 McDiarmid argument for randomly perturbed graphs
To prove (1), it suffices to show that the inequality holds when conditioning each side on each of (a), (b) and (c). This holds with equality if (a) or (b) holds: if (a) holds, both sides are 11, and if (b) holds, both sides are 0. Now consider (c).
@@ 3 Almost spanning rainbow trees in random graphs
Observe that $G_{p_{0}}$ contains a copy of $T$ with high probability, by Proposition $3.$ Let $\mathbf{G}(n,p_{0})$ be the binomial random graph on $V$, coupled with $G_{p_{0}}$ so that $G_{p_{0}}\subseteq\mathbf{G}(n,p_{0})$. Then it follows that with high probability $\mathbf{G}(n,p_{0})$ contains a graph in $\mathcal{H}$.
@@
Let $G_{0}$ be a graph on vertex set $[n]$ with minimum degree at least $\delta n$. Let $G\sim G_{0}\cup\mathbf{G}(n,C/n)$ and suppose $G$ is uniformly coloured in [n1][n-1]. Let $T$ be an $n$-vertex tree with maximum degree at most $d$ that we wish to embed in a rainbow fashion in $G$. To aid our embedding, we seek simple structures in $T$, for which we use the following observation of Krivelevich [23], where we recall that a bare path in $T$ is a path whose interior vertices have degree 22 in $T$.
@@ 4.1 Embedding trees with many leaves
We first embed $T\setminus L$ in a rainbow fashion in $\mathbf{G}(n,C/n)$ using Theorem 1.4. Completing this to a rainbow embedding of $T$ essentially amounts to finding a rainbow perfect matching between the image of $M$ and the uncovered vertices of $V$, which uses all the unused colours. This can be reduced to finding a rainbow directed Hamilton cycle in a suitable auxiliary edge-coloured perturbed directed graph. For that we will apply the following result of the authors, where we recall that $\mathbf{D}(n,p)$ denotes the binomial random digraph on $n$ vertices with edge probability pp.
@@ 4.2 Embedding trees with many bare paths
We will state here several lemmas (namely Lemmas 4.6, 4.8 and 4.9) from a manuscript [21] by the first two authors, where they proved an undirected version of Theorem 4.5. These lemmas all have analogues in [22] for the directed setting, and in most cases have very similar proofs, but the directed versions do not immediately imply the undirected ones (due to parallel directed edges xyxy and yxyx being coloured independently).
@@ 6 Embedding trees with many bare paths
Recall that for each pair $(x_{i},y_{i})$ we built $80k$ gadgets and $80k+1$ connecting paths. As we will see shortly, when using the absorbing structure, for each pair we will use the absorbing paths (each of length $10$, c.f. Figure 1) in precisely $2k$ gadgets, and the avoiding path (each of length 99, c.f. Figure 1) in every other gadget. Together with the connecting paths (each of length 33), we will get a path of length
@@
  • each path has length 77 and all its 66 internal vertices in $V_{\operatorname{flex}}$;
@@
and observe that $P_{i}$ is a path with endpoints $x_{i}$ and $y_{i}$. Each $x\in R_{i}$ lies in exactly one edge of $M_{1}$ and exactly one edge of $M_{2}$, so $P_{i}$ consists of $2|R_{i}|=2k$ absorbing paths (of length $10$), 80k2k80k-2k avoiding paths (of length 99), and $80k+1$ connecting paths (of length 33). Hence, $P_{i}$ has length $\ell$ (c.f. Equation 4). Finally define
@@ 7 Concluding remarks
    For any disjoint vertex subsets AA and $B$ with $\left|A\right|=a$, $\left|B\right|=b$ and $abp\geq 250n$, the number of edges between them is at least $abp/4$ and at most $3abp/2$.
@@
    Every subset of vertices AA of size $a\leq n/4$ spans at most $apn/2$ edges.
@@
Finally we consider the case $t\geq\frac{\log D}{D}\cdot n$. Set $Z:=U\setminus(X\cup Y)$, and notice that $e_{G_{p}}(X,Z)=0$, since $G^{\ast}[U]$ is an induced subgraph of $G_{p}$. From $\left|U\right|\geq(2d+2)t$, $|X|=t$ and $\left|Y\right|\leq(d+1)t$ it follows that $\left|Z\right|\geq dt$. Let $B_{t}$ be the event that there exist vertex subsets of size $t$ and dtdt, with no edge between them in $G_{p}$. Then

### v3

@@ 1 Introduction
Given $\delta\in(0,1)$, we define $\mathcal{G}_{\delta,n}$ to be the family of graphs on vertex set $[n]$ with minimum degree at least $\delta n$, and we let $\mathbf{G}(n,p)$ be the binomial random graph on vertex set $[n]$ with edge probability $p$. One of the central themes in extremal combinatorics is determining the minimum degree threshold for a given graph property $\mathcal{P}$, i.e. how large $\delta$ needs to be so that every $G\in\mathcal{G}_{\delta,n}$ satisfies $\mathcal{P}$. Similarly, probabilistic combinatorics aims to determine how large $p$ needs to be for $\mathbf{G}(n,p)$ to satisfy $\mathcal{P}$ with high probability [^1] Formally, we say that a sequence of events $(A_{n})_{n\in\mathbb{N}}$ holds with high probability if $\mathbb{P}[A_{n}]\rightarrow 1$ as $n\rightarrow\infty$. . Bohman, Frieze and Martin [8] provided a connection between the extremal and the random graph settings by introducing the randomly perturbed graph model. For a given $\delta\in(0,1)$, this is defined as $G_{0}\cup\mathbf{G}(n,p)$ where $G_{0}\in\mathcal{G}_{\delta,n}$, i.e. as the graph on $[n]$ whose edge set is the union of the edges of a deterministic graph $G_{0}$ with minimum degree at least $\delta n$ and the edges of a random graph $\mathbf{G}(n,p)$ on the same vertex set. For a given $\delta$ and a given graph property $\mathcal{P}$, a pivotal question in the area is to determine how large $p$ needs to be so that for every $G_{0}\in\mathcal{G}_{\delta,n}$, with high probability, $G_{0}\cup\mathbf{G}(n,p)$ satisfies $\mathcal{P}$. More precisely, we say that $\hat{p}=\hat{p}(\delta,\mathcal{P},n)$ is a perturbed threshold for the property $\mathcal{P}$ at $\delta$ if there are constants $C>c>0$ such that for any $p\geq C\hat{p}$ and for any sequence of $n$-vertex graphs $(G_{n})_{n\in\mathbb{N}}$ with $G_{n}\in\mathcal{G}_{\delta,n}$ we have $\lim_{n\to\infty}\mathbb{P}\big(G_{n}\cup G(n,p)\in\mathcal{P}\big)=1$, and for any $p\leq c\hat{p}$ there exists a sequence of $n$-vertex graphs $(G_{n})_{n\in\mathbb{N}}$ with $G_{n}\in\mathcal{G}_{\delta,n}$ such that $\lim_{n\to\infty}\mathbb{P}\big(G_{n}\cup G(n,p)\in\mathcal{P}\big)=0$.
@@
Given a graph $G$, a vertex $v\in V(G)$ and a subset $X\subseteq V(G)$, $E_{G}(v,X)$ denotes the set of edges of the form $vx$ with $x\in X$, and $N_{G}(X)$ denotes the set of vertices with at least one neighbour in $X$. A bare path in $G$ is a path whose interior vertices have degree $2$ in $G$. For a graph $G$ and $p\in[0,1]$, the $p$-random subgraph of $G$, denoted by $G_{p}$, is the random graph resulting from sampling each edge of $G$ independently with probability $p$.
A digraph $D$ is a set of vertices together with a set of ordered pairs of distinct vertices and the minimum semi-degree $\delta^{0}(D)$ of $D$ is the minimum over in- and out-degrees of vertices in $D$. Moreover $\mathbf{D}(n,p)$ denotes the binomial random digraph on $n$ vertices, that is the digraph on $[n]$ where each ordered pair of distinct vertices forms a directed edge independently with probability $p$.
@@ 2 McDiarmid argument for randomly perturbed graphs
To prove (1), it suffices to show that the inequality holds when conditioning each side on each of (a), (b) and (c). This holds with equality if (a) or (b) holds: if (a) holds, both sides are $1$, and if (b) holds, both sides are 0. Now consider (c).
@@ 3 Almost spanning rainbow trees in random graphs
Observe that $G_{p_{0}}$ contains a copy of $T$ with high probability, by Proposition 3.3. Let $\mathbf{G}(n,p_{0})$ be the binomial random graph on $V$, coupled with $G_{p_{0}}$ so that $G_{p_{0}}\subseteq\mathbf{G}(n,p_{0})$. Then it follows that with high probability $\mathbf{G}(n,p_{0})$ contains a graph in $\mathcal{H}$.
@@
Let $G_{0}$ be a graph on vertex set $[n]$ with minimum degree at least $\delta n$. Let $G\sim G_{0}\cup\mathbf{G}(n,C/n)$ and suppose $G$ is uniformly coloured in [n1][n-1]. Let $T$ be an $n$-vertex tree with maximum degree at most $d$ that we wish to embed in a rainbow fashion in $G$. To aid our embedding, we seek simple structures in $T$, for which we use the following observation of Krivelevich [23], where we recall that a bare path in $T$ is a path whose interior vertices have degree $2$ in $T$.
@@ 4.1 Embedding trees with many leaves
We first embed $T\setminus L$ in a rainbow fashion in $\mathbf{G}(n,C/n)$ using Theorem 1.4. Completing this to a rainbow embedding of $T$ essentially amounts to finding a rainbow perfect matching between the image of $M$ and the uncovered vertices of $V$, which uses all the unused colours. This can be reduced to finding a rainbow directed Hamilton cycle in a suitable auxiliary edge-coloured perturbed directed graph. For that we will apply the following result of the authors, where we recall that $\mathbf{D}(n,p)$ denotes the binomial random digraph on $n$ vertices with edge probability $p$.
@@ 4.2 Embedding trees with many bare paths
We will state here several lemmas (namely Lemmas 4.6, 4.8 and 4.9) from a manuscript [21] by the first two authors, where they proved an undirected version of Theorem 4.5. These lemmas all have analogues in [22] for the directed setting, and in most cases have very similar proofs, but the directed versions do not immediately imply the undirected ones (due to parallel directed edges $xy$ and $yx$ being coloured independently).
@@ 6 Embedding trees with many bare paths
Recall that for each pair $(x_{i},y_{i})$ we built $80k$ gadgets and $80k+1$ connecting paths. As we will see shortly, when using the absorbing structure, for each pair we will use the absorbing paths (each of length $10$, c.f. Figure 1) in precisely $2k$ gadgets, and the avoiding path (each of length $9$, c.f. Figure 1) in every other gadget. Together with the connecting paths (each of length $3$), we will get a path of length
@@
  • each path has length $7$ and all its $6$ internal vertices in $V_{\operatorname{flex}}$;
@@
and observe that $P_{i}$ is a path with endpoints $x_{i}$ and $y_{i}$. Each $x\in R_{i}$ lies in exactly one edge of $M_{1}$ and exactly one edge of $M_{2}$, so $P_{i}$ consists of $2|R_{i}|=2k$ absorbing paths (of length $10$), 80k2k80k-2k avoiding paths (of length $9$), and $80k+1$ connecting paths (of length $3$). Hence, $P_{i}$ has length $\ell$ (c.f. Equation 4). Finally define
@@ 7 Concluding remarks
    For any disjoint vertex subsets $A$ and $B$ with $\left|A\right|=a$, $\left|B\right|=b$ and $abp\geq 250n$, the number of edges between them is at least $abp/4$ and at most $3abp/2$.
@@
    Every subset of vertices $A$ of size $a\leq n/4$ spans at most $apn/2$ edges.
@@
Finally we consider the case $t\geq\frac{\log D}{D}\cdot n$. Set $Z:=U\setminus(X\cup Y)$, and notice that $e_{G_{p}}(X,Z)=0$, since $G^{\ast}[U]$ is an induced subgraph of $G_{p}$. From $\left|U\right|\geq(2d+2)t$, $|X|=t$ and $\left|Y\right|\leq(d+1)t$ it follows that $\left|Z\right|\geq dt$. Let $B_{t}$ be the event that there exist vertex subsets of size $t$ and $dt$, with no edge between them in $G_{p}$. Then